A battery has emf 13.2V and internal resistance
24mΩ
24mΩ
. If the load current is 20.0A, find the terminal voltage o f the battery
12.7V
14.5V
16.8V
17.7V
1
Expert's answer
2016-04-25T08:46:06-0400

V = emf - I * r

V = 13.2 - 20.0 * 24*10^-3 = 12.72 V